The smell of a burning match is due to the formation of sulfur oxides. The relevant reaction is S + O2 → SO2. If 4.8g of sulfur is burned in this manner, what mass of sulfur dioxide is produced?


1
Expert's answer
2021-09-10T00:41:16-0400

Mole ratio of S : SO2 = 1:1

Molar mass of sulfur = 32.065

= 4.8/32.065 = 0.1497

Molar mass of SO2 = 64.066 g

= 0.1497× 64.066 = 9.59g